# Josephson traveling-wave parametric amplifier based on low-intrinsic-loss coplanar lumped-element waveguide

### v2

@@ IV.2 Noise performance and saturation
Next, we characterize the noise performance of the Tukey CP-JTWPA with the configuration in Fig. 4(c) and microwave network described in Fig. 9. This requires precise knowledge of power levels at the CP-JTWPA input, which demands accurate calibration of the system gain from room temperature down to the device output port. We performed this calibration at 5.563 GHz using a waveguide-QED approach with a qubit-coupled transmission line as a reference device [42]. By analyzing the frequency- and power-dependent transmission of this qubit-coupled line, we established the system input attenuation $A_{\mathrm{sys}}$ between room temperature and the qubit. To extend this calibration to the CP-JTWPA measurements, we employed two microwave switches to alternately connect either the CP-JTWPA or the qubit-coupled line to identical measurement paths, ensuring equivalent signal paths through cables of matched lengths (see Appendix F for details of the noise calibration setup and procedures). This configuration allowed us to translate the calibrated attenuation from the qubit reference to the CP-JTWPA. The system gain through the CP-JTWPA was then determined by measuring the overall transmission and compensating for the calibrated input attenuation.
Figure 6: Saturation power of the Tukey CP-JTWPA measured at 20.5 dB gain and 5.563 GHz. The CP-JTWPA exhibits a good gain linearity from $-150$ dBm to $-104$ dBm in input signal power, and saturates at $-99$ dBm.
With the calibrated system gain, we measured noise spectra at room temperature with the CP-JTWPA pump turned on and off to determine both the added noise by the CP-JTWPA, $N_{\mathrm{JTWPA}}$, and the system noise $N_{\mathrm{sys}}$ referred to the CP-JTWPA input port. As shown in Fig. 5, we achieved near-quantum-limited performance at the calibrated frequency with an added noise of 0.63 quanta for gains up to 21 dB. At this gain, the system noise temperature decreased to 320 mK (1.2 quanta), relatively low comparing to most previous implementations [9, 11, 14].
Using the same calibration, we also extracted the saturation power $P_{\mathrm{1dB}}$ by performing a probe power sweep and observing at what power $G_{\mathrm{JTWPA}}$ compresses by 1 dB, shown in Fig. 6. With a reference gain of 20.5 dB, the CP-JTWPA exhibits excellent linearity in gain throughout a wide range of input signal power from $-150$ dBm to $-104$ dBm. Saturation was measured at $-99$ dBm.
@@ V Conclusion and discussion
Through systematic comparison of window functions, we established that the Tukey window achieves a good balance between phase-matching strength and ripple suppression. The resulting performance, with over 20-dB gain across a 5-GHz bandwidth with ideal impedance matching while maintaining intrinsic gain variations within ±$\pm 2$ dB, demonstrated the effectiveness of this approach. In a more realistic impedance-mismatched environment, our Tukey device still demonstrated 17-20 dB gain over 4.8 GHz, with added noise below 0.63 quanta and a saturation power of $-99$ dBm.
@@ ACKNOWLEDGEMENT
Appendix A DEVICE PARAMETERS AND MEASUREMENT CONFIGURATIONS
@@
Number of cells (cells) $N$ 1600 2400 2397
Cell separation ($\mu$m) $a$ 20 20 20
Stub length [^4] Values shown refers roughly to the mean of the modulated parameter arrays. ($\mu$m) $l_{\text{stub}}$ 80 100 100
@@
[…]
@@
Pump power [^5] The pump power is estimated by the power applied in fitting the theoretical gain prediction to the gain curve together with other parameters of the CP-JTWPAs. (dBm) $P_{\mathrm{p}}$ $-68.8$ $-71.4$ $-70.3$
@@
The devices differ primarily in their modulation depth, window functions and number of unit cells, while sharing similar basic architecture and fabrication recipes. Due to the challenges in direct parameter measurements of JTWPAs, the circuit parameters were determined through a combination of design considerations and experimental validation. The stub capacitance was initially estimated from finite-element electromagnetic simulations, the junction inductance from normal-state resistance measurements, and the junction capacitance from device geometry. These parameters were then refined through fitting simulated transmission and gain characteristics to experimental measurements.
To achieve sinusoidal impedance modulation with amplitude $m$, we coordinate variations in both junction area and stub length. Since the characteristic impedance $Z=\sqrt{(L_{\mathrm{J}}+L_{\mathrm{s}})/C_{\mathrm{g}}}$ includes the nearly constant geometric inductance $L_{\mathrm{s}}$, we first implement a sinusoidal modulation of junction area to vary $L_{\mathrm{J}}$. We then calculate the corresponding stub length modulation to adjust $C_{\mathrm{g}}$ accordingly, ensuring the overall impedance profile follows the intended sinusoidal pattern with amplitude $m$.
Throughout our device development cycle, we progressively increased isolation attenuation from 3-dB (boxcar) to 6-dB (Hann) and finally 10-dB (Tukey) as we better understood the impact of impedance matching on device characterization. While these differences might suggest an unfair comparison, simulations fully support the observed gain ripples in the boxcar device. Furthermore, even when tested without output impedance matching, the Tukey device still exhibited significantly fewer ripples than the boxcar design. These consistent results across different measurement conditions confirm the inherent effectiveness of our windowed modulation approach.
@@ Appendix B COUPLED-MODE ANALYSIS OF WINDOWED PERIODIC MODULATION
where $z$ is the position along the transmission line $(0\leq z\leq l)$, $m$ is the modulation depth $(0<m<1)$, $\beta_{\mathrm{B}}=\pi/\Lambda$ is the Bragg wave vector with $\Lambda$ being the period of the modulation, and $W(z)$ is the window function. The resulting modulated impedance can be expressed as
@@
where $\kappa=\beta_{0}m/4$ is the coupling coefficient between the forward and backward propagating waves, and $\Delta_{\beta}=\beta_{0}-\beta_{\mathrm{B}}$ is the detuning parameter. To obtain physically intuitive solutions that reveal the roles of window function and coupling strength, we employ a transfer-matrix formalism, which connects the voltage wave amplitudes at different points in a passive medium. Given the linear nature of Eqs. (B.11) and (B.12), we expect the field amplitudes at any position $z$ to be linearly related to those at position $z_{0}$ through a 2$\times$2 transfer matrix $T(z_{0},z)$,
@@
$\begin{split}&\begin{pmatrix}0&j\kappa W(z)e^{j2\Delta_{\beta}z}\\ -j\kappa W(z)e^{-j2\Delta_{\beta}z}&0\end{pmatrix}T(z_{0},z)\begin{pmatrix}V^{+}(z_{0})\\ V^{-}(z_{0})\end{pmatrix}\\ &=\frac{d}{dz}[T(z_{0},z)]\begin{pmatrix}V^{+}(z_{0})\\ V^{-}(z_{0})\end{pmatrix}.\end{split}$ (B.14)
@@
[…]
@@
$[A(z^{\prime}),A(z^{\prime\prime})]=2j\kappa^{2}W(z^{\prime})W(z^{\prime\prime})\begin{pmatrix}\sin\phi&0\\ 0&-\sin\phi\end{pmatrix},$ (B.20)
@@
[…]
@@ C.2 Example parameters for comparing analytical and numerical calculations
[…]
@@
Appendix D HARMONIC-BALANCE SIMULATIONS OF CP-JTWPAS
@@
Figure 7: Harmonic balance simulations of the CP-JTWPA with different window functions. (a-c) Simulated scattering parameters for the three CP-JTWPA designs: (a) boxcar window (device A), (b) Hann window (device B), and (c) Tukey window (device C). All S-parameters ($|S_{11}|^{2}$, $|S_{12}|^{2}$, $|S_{21}|^{2}$, and $|S_{22}|^{2}$) are shown over the 4–12 GHz frequency range with configuration given in Appendix A. The high-gain regions in $|S_{21}|^{2}$ (green) correspond to the phase-matched parametric amplification bands, while the bandgap regions appear at approximately 7.9 GHz, 8.1 GHz, and 8.8 GHz for the three devices, respectively. (d-f) Corresponding quantum efficiency ratios (QE/$\mathrm{QE_{ideal}}$) for each device, showing the deviation from ideal JTWPA behavior. The ratio approaches unity in most of the operating bandwidth but drops significantly within and near the bandgap frequencies.
@@
Harmonic-balance simulations for our CP-JTWPAs were performed using the library called JosephsonCircuits [41], based on the parameters presented in Appendix A, assuming perfect impedance matching at both ends. We swept with 10-MHz steps over the 4–12-GHz range to capture potential ripples in the S-parameters. To obtain sufficient accuracy in quantum efficiency estimations, we used the harmonic-balance method with 20 harmonics of the pump frequency $\omega_{\mathrm{p}}$ to solve the nonlinear circuit response, and 10 harmonics in the signal-idler space for the linearized small-signal analysis. In our 4WM implementation, we included pump harmonics $n\omega_{\mathrm{p}}$ for $n=1,3,5,..,19$ with odd harmonics dominating the mixing process, and signal/idler mixing products of the form $\omega_{\mathrm{s,i}}\pm 2n\omega_{\mathrm{p}}$ for $n=0,1,2,3,4$. This approach captured 4WM processes up to the 8th-order sideband, with the highest-frequency side bands showing gain as low as $-80$ dB compared to the input signal power, sufficiently negligible for accurate simulation results. From our simulation, we observed that the quantum efficiency (QE) can reach over 90% of the ideal QE given by [44]
$QE_{\mathrm{ideal}}(G)=\frac{1}{2-1/G}.$ (D.1)
@@
Shown in Fig. 7, the $|S_{21}|^{2}$ parameters from these simulations were used for fitting experimental data in the main text. The simulations predict significant reflected gain ($|S_{11}|^{2}$ and $|S_{22}|^{2}$) near the bandgap in the boxcar device, while reflections remain low ($<$10 dB) for the Hann and Tukey devices due to the suppression of reflections around the bandgap by the windowed modulations. The backward gain $|S_{12}|^{2}$ generally remains below 0 dB in all simulations, with a small frequency range showing approximately 1 dB gain over the low-frequency region in the Tukey device with this pump configuration. The mechanism behind negative backward gain regions remains unclear. It is worth further investigations as it potentially allows a non-reciprocal amplification.
All devices showed relatively high quantum efficiency in simulations. This can be attributed to the use of relatively weak pump power compared to the junction critical current, which reduced interactions between the signal and higher-frequency modes. The Hann and Tukey devices demonstrate higher overall quantum efficiency for two reasons. First, reflections around the bandgap rapidly deteriorate quantum efficiency in the boxcar device. Second, the additional parallel capacitance from larger $C_{\mathrm{g}}$ in the Hann and Tukey devices creates stronger curvature in their dispersion relations, further suppressing interactions between the signal mode and higher-frequency modes.
@@
It should be noted that for the Tukey window case, we encountered convergence challenges in the simulation, particularly near the bandgap region. The sharp ripples visible in these regions in other S-parameters than $|S_{21}|^{2}$ may therefore be numerical artifacts rather than representing the actual device response. This convergence issue likely stems from two factors: first, the shorter taper length in the Tukey window compared to the Hann window may retain some abrupt transitions in the impedance profile; second, we positioned the pump frequency very close to the bandgap where residual reflection ripples still exist. While residual ripples were also present in the boxcar case, we did not encounter similar convergence issues because the pump was necessarily placed further from the bandgap due to the slow decay of reflections in that design.

### v3

@@ IV.2 Noise performance and saturation
Next, we characterize the noise performance of the Tukey CP-JTWPA with the configuration in Fig. 4(c) and microwave network described in Fig. 10. This requires precise knowledge of power levels at the CP-JTWPA input, which demands accurate calibration of the system gain from room temperature down to the device output port. We performed this calibration at 5.563 GHz using a waveguide-QED approach with a qubit-coupled transmission line as a reference device [44]. By analyzing the frequency- and power-dependent transmission of this qubit-coupled line, we established the system input attenuation $A_{\mathrm{sys}}$ between room temperature and the qubit. To extend this calibration to the CP-JTWPA measurements, we employed two microwave switches to alternately connect either the CP-JTWPA or the qubit-coupled line to identical measurement paths, ensuring equivalent signal paths through cables of matched lengths (see Appendix G for details of the noise calibration setup and procedures). This configuration allowed us to translate the calibrated attenuation from the qubit reference to the CP-JTWPA. The system gain through the CP-JTWPA was then determined by measuring the overall transmission and compensating for the calibrated input attenuation.
Figure 6: Saturation power of the Tukey CP-JTWPA measured at 20.5 dB gain and 5.563 GHz. The CP-JTWPA exhibits a good gain linearity from 150-150 dBm to 104-104 dBm in input signal power, and saturates at 99-99 dBm.
With the calibrated system gain, we measured noise spectra at room temperature with the CP-JTWPA pump turned on and off to determine both the added noise by the CP-JTWPA, $N_{\mathrm{JTWPA}}$, and the system noise $N_{\mathrm{sys}}$ referred to the CP-JTWPA input port. As shown in Fig. 5, we achieved near-quantum-limited performance at the calibrated frequency with an added noise of 0.68 quanta for gains up to 21 dB, corresponding to a quantum efficiency at $84\%$ of an ideal quantum-noise-limited amplifier (see Appendix D for details). With this gain configuration, the input-referred system noise temperature decreased to 382 mK (1.43 quanta), relatively low comparing to most previous implementations [9, 11, 14].
Using the same calibration, we also extracted the saturation power $P_{\mathrm{1dB}}$ by performing a probe power sweep and observing at what power $G_{\mathrm{JTWPA}}$ compresses by 1 dB, shown in Fig. 6. With a reference gain of 20.5 dB, the CP-JTWPA exhibits excellent linearity in gain throughout a wide range of input signal power from 150-150 dBm to 104-104 dBm. Saturation was measured at 99-99 dBm.
@@ V Conclusion and discussion
Through systematic comparison of window functions, we established that the Tukey window achieves a good balance between phase-matching strength and ripple suppression. The resulting performance, with over 20-dB gain across a 5-GHz bandwidth with ideal impedance matching while maintaining intrinsic gain variations within $\pm 2$ dB, demonstrated the effectiveness of this approach. In a more realistic impedance-mismatched environment, our Tukey device still demonstrated 17-20 dB gain over 4.8 GHz, with added noise below 0.63 quanta and a saturation power of 99-99 dBm.
@@ ACKNOWLEDGEMENT
Appendix A DEVICE PARAMETERS AND MODULATION CONFIGURATIONS
@@
Number of cells (cells) NN 2400 2400 2397
[…]
@@
Junc. capacitance [^1] This value represents the intrinsic junction capacitance $C_{\mathrm{J}}$ estimated from simulation parameters. It is calculated by subtracting the series capacitance contribution from adjacent open stubs (approximately $C_{\mathrm{g}}/2$) from the total parallel capacitance: $C_{\mathrm{J}}\approx C_{\mathrm{parallel}}-C_{\mathrm{g}}/2$. [^4] Values shown refers roughly to the mean of the modulated parameter arrays. (fF) $C_{\mathrm{J}}$ 48.5 54 54.5
@@
Pump power [^5] The pump power is estimated by the power applied in fitting the theoretical gain prediction to the gain curve together with other parameters of the CP-JTWPAs. (dBm) $P_{\mathrm{p}}$ 68.8-68.8 71.4-71.4 70.3-70.3
@@
[…]
To achieve sinusoidal impedance modulation with amplitude $m$, we coordinate variations in both junction area and stub length. The characteristic impedance is given by $Z=\sqrt{(L_{\mathrm{J}}+L_{\mathrm{s}})/C_{\mathrm{g}}}$, where $L_{\mathrm{s}}$ is almost constant over a wide range of stub length. To achieve an 8% impedance modulation through stub length alone would require roughly 16% variation in $C_{\mathrm{g}}$, leading to large geometric changes. Instead, we apply a correlated modulation on both junction area (varying $L_{\mathrm{J}}$) and stub length (varying $C_{\mathrm{g}}$), such that their combined effect produces the target impedance profile $Z(z)=Z_{0}[1+m\cdot W(z)\sin(2\pi z/\Lambda)]$. This reduces the required variation in the stub length to approximately 8% while maintaining the same overall impedance-modulation amplitude.
@@ Appendix B COUPLED-MODE ANALYSIS OF WINDOWED PERIODIC MODULATION
where $z$ is the position along the transmission line $(0\leq z\leq l)$, mm is the modulation depth $(0<m<1)$, $\beta_{\mathrm{B}}=\pi/\Lambda$ is the Bragg wave vector with $\Lambda$ being the period of the modulation, and $W(z)$ is the window function. The resulting modulated impedance can be expressed as
@@
where $\kappa=\beta_{0}m/4$ is the coupling coefficient between the forward and backward propagating waves, and $\Delta_{\beta}=\beta_{0}-\beta_{\mathrm{B}}$ is the detuning parameter. To obtain physically intuitive solutions that reveal the roles of window function and coupling strength, we employ a transfer-matrix formalism, which connects the voltage wave amplitudes at different points in a passive medium. Given the linear nature of Eqs. (B.11) and (B.12), we expect the field amplitudes at any position $z$ to be linearly related to those at position $z_{0}$ through a $\times2$ transfer matrix $T(z_{0},z)$,
@@
[…]
@@
$\Omega_{1}(z_{0},z)$ $=\int_{z_{0}}^{z}A(z^{\prime})dz^{\prime},$ (B.18)
$\Omega_{2}(z_{0},z)$ $=\frac{1}{2}\int_{z_{0}}^{z}\int_{z_{0}}^{z^{\prime}}[A(z^{\prime}),A(z^{\prime\prime})]dz^{\prime\prime}dz^{\prime}.$ (B.19)
@@
$A(z^{\prime}),A(z^{\prime\prime})]=2j\kappa^{2}W(z^{\prime})W(z^{\prime\prime})\begin{pmatrix}\sin\phi&0\\ 0&-\sin\phi\end{pmatrix}$, (B.20)
@@
$F_{\mathrm{Hann}}(0,l)$ $=\int_{0}^{l}W_{\mathrm{Hann}}(z)e^{j2\Delta_{\beta}z}dz$
$=\frac{1}{2}\int_{0}^{l}e^{j2\Delta_{\beta}z}dz-\frac{1}{2}\int_{0}^{l}\cos(2\pi z/l)e^{j2\Delta_{\beta}z}dz$
@@ C.2 Example parameters for comparing analytical and numerical calculations
$L$ $=5.3\text{ pH/}\mu\text{m}$ (series inductance per unit length)
$C$ $=1.9\text{ fF/}\mu\text{m}$ (shunt capacitance per unit length)
@@
Appendix D GAIN SIMULATIONS AND QUANTUM EFFICIENCY ANALYSIS OF CP-JTWPAS
@@
Figure 7: Harmonic balance simulations of the CP-JTWPA with different window functions. (a-c) Simulated scattering parameters for the three CP-JTWPA designs: (a) boxcar window (device A), (b) Hann window (device B), and (c) Tukey window (device C). All S-parameters ($|S_{11}|^{2}$, $|S_{12}|^{2}$, $|S_{21}|^{2}$, and $|S_{22}|^{2}$) are shown over the 4–12 GHz frequency range with configuration given in Appendix A. The high-gain regions in $|S_{21}|^{2}$ (green) correspond to the phase-matched parametric amplification bands, while the band-gap regions appear at approximately 7.9 GHz, 8.1 GHz, and 8.8 GHz for the three devices, respectively. (d-f) Corresponding quantum efficiency ratios ($\eta/\eta_{\mathrm{ideal}}$) for each device, showing the deviation from ideal JTWPA behavior. The ratio approaches unity in most of the operating bandwidth but drops significantly within and near the band-gap frequencies.
D.1 Gain simulations
@@
We simulate the gain of our CP-JTWPAs using harmonic-balance calculations with the JosephsonCircuits.jl library [43] and the parameters given in Appendix A, assuming perfect impedance matching at both ends. We swept with 10-MHz steps over the 4–12-GHz range to capture potential ripples in the S-parameters. To obtain sufficient accuracy in our nonlinear analysis, we used the harmonic-balance method with 20 harmonics of the pump frequency $\omega_{\mathrm{p}}$ to solve the nonlinear circuit response, and 10 harmonics in the signal-idler space for the linearized small-signal analysis. In our 4WM implementation, we included pump harmonics $n\omega_{\mathrm{p}}$ for $n=1,3,5,..,19$ with odd harmonics dominating the mixing process, and signal/idler mixing products of the form $\omega_{\mathrm{s,i}}\pm 2n\omega_{\mathrm{p}}$ for $n=0,1,2,3,4$. This approach captured 4WM processes up to the 8th-order sideband, with the highest-frequency side bands showing gain as low as 80-80 dB compared to the input signal power, sufficiently negligible for accurate simulation results.
@@
Shown in Fig. 7, the $|S_{21}|^{2}$ parameters from these simulations were used for fitting experimental data in the main text. The simulations predict significant reflected gain ($|S_{11}|^{2}$ and $|S_{22}|^{2}$) near the band gap in the boxcar device, while reflections remain low (<<10 dB) for the Hann and Tukey devices due to the suppression of reflections around the band gap by the windowed modulations. The backward gain $|S_{12}|^{2}$ generally remains below 0 dB in all simulations, with a small frequency range showing approximately 1 dB gain over the low-frequency region in the Tukey device with this pump configuration. The mechanism behind negative backward gain regions remains unclear. It is worth further investigations as it potentially allows a non-reciprocal amplification.
@@
It should be noted that for the Tukey window case, we encountered convergence challenges in the simulation, particularly near the band-gap region. The sharp ripples visible in these regions in other S-parameters than $|S_{21}|^{2}$ may therefore be numerical artifacts rather than representing the actual device response. This convergence issue likely stems from two factors: first, the shorter taper length in the Tukey window compared to the Hann window may retain some abrupt transitions in the impedance profile; second, we positioned the pump frequency very close to the band gap where residual reflection ripples still exist. While residual ripples were also present in the boxcar case, we did not encounter similar convergence issues because the pump was necessarily placed further from the band-gap due to the slow decay of reflections in that design.
D.2 Quantum efficiency analysis
Regarding quantum efficiency $\eta$, our simulations predict that the amplifiers can theoretically reach over 90% of the ideal quantum efficiency $\eta_{\mathrm{ideal}}$. This high performance can be attributed to the use of relatively weak pump power compared to the junction critical current, which reduced interactions between the signal and higher-frequency modes. The Hann and Tukey devices demonstrate higher overall $\eta$ than the boxcar device for two reasons. First, reflections around the band gap rapidly deteriorate $\eta$ in the boxcar device. Second, the additional parallel capacitance from larger $C_{\mathrm{g}}$ in the Hann and Tukey devices creates stronger curvature in their dispersion relations, further suppressing interactions between the signal mode and higher-frequency modes.
In experiments, with the measured gain $G$ and added noise $A$ of an amplifier, its $\eta$ can be calculated as [46]
$\eta(A)=\frac{1}{1+2A},$ (D.1)
and it can be compared to that of an ideal quantum-noise-limited amplifier, given by
$\eta_{\mathrm{ideal}}(G)=\frac{1}{2-1/G}.$ (D.2)
The ratio $\eta/\eta_{\mathrm{ideal}}$ can serve as a figure of merit for near-quantum-noise-limited amplifiers.
Appendix E EFFECT OF IMPEDANCE MISMATCH ON JTWPA AND THE USE OF ATTENUATORS
Figure 8: Harmonic-balance simulations of the boxcar CP-JTWPA under different impedance matching and attenuation conditions. (a,b) Baseline-transmission characteristics with pump off (pump power below 110-110 dBm): (a) ideal 50-$\Omega$ port impedance and (b) mismatched 40-$\Omega$ port impedance yielding $|S_{11}|^{2}=-15$ dB reflection. (c–f) Gain profiles with the CP-JTWPA sandwiched between attenuators each providing: (c) 0 dB, (d) 3 dB, (e) 6 dB, and (f) 10 dB attenuation.
Throughout this work, when characterizing the intrinsic gain of the CP-JTWPAs, cryogenic attenuators were used to sandwich the JTWPA packages. These attenuators serve to isolate the JTWPAs from external impedance mismatches, with attenuation values chosen to minimize signal-to-noise ratio degradation while providing adequate isolation, allowing the measurements to accurately reflect the gain profiles of the individual designs. We progressively increased the installed attenuation from 3 dB (boxcar) to 6 dB (Hann) and finally 10 dB (Tukey), as we observed higher gains from successive samples and recognized that stronger attenuation was needed. While these different attenuation values might initially appear to compromise the fairness of device comparisons, the following analysis demonstrates that the observed ripple characteristics represent intrinsic device behavior rather than measurement artifacts.
E.1 Origin of gain-profile ripples and attenuation required for their suppression
The role of attenuators is to suppress external reflections that could create measurement artifacts, allowing us to observe the device’s intrinsic gain behavior. An $N$-dB attenuator attached to one port of the CP-JTWPA provides a minimum return loss of 2N2N\! dB, since signals must traverse the attenuator twice when reflected by external components (assuming the attenuator itself has excellent return loss, typically >>25 dB). This means that a device sandwiched between two $N$-dB attenuators experiences a minimum round-trip loss of 4N4N\! dB from the attenuators alone. When combined with the return losses $R$ from external components (cables and circulators typically providing 15–20 dB each), the total round-trip loss becomes 4N+2R4N\!+2R\! dB.
Gain ripples arise when impedance mismatches at the JTWPA boundaries cause signals to undergo multiple round trips, with each forward pass through the device providing additional amplification. In our near-lossless devices, if the total round-trip loss approaches or falls below the single-pass gain, signals can bounce repeatedly between the input and output terminations, accumulating amplified power over multiple traversals. From the receiver’s perspective, this results in enhanced power accumulation over the measurement integration window, manifesting as gain ripples when sweeping across frequency.
For our boxcar device sandwiched between 3-dB attenuators, the minimum round-trip loss from the attenuators alone is 12 dB. Adding the return losses from external components, the total round-trip loss typically exceeds 40 dB, significantly exceeding the 10–13-dB gain of our boxcar CP-JTWPA, which should adequately suppress ripples from external reflections. The observation of pronounced gain ripples despite this suppression provides strong evidence for significant internal reflections within the device itself, consistent with our theoretical predictions for periodic-modulation effects.
E.2 Simulated gain profiles of boxcar CP-JTWPA under impedance mismatch
To validate the attenuation levels needed for the boxcar device to observe its intrinsic gain profile, we simulated its baseline transmission and gain under different impedance-matching and pump configurations as shown in Fig. 8. Accurately modeling frequency-dependent impedance mismatches in components connecting to JTWPAs requires precise knowledge of their scattering parameters at cryogenic temperatures, which are difficult to obtain and demand extensive specialized measurements. Instead, we model impedance mismatch in our simulations by assuming a constant port impedance of 40 $\Omega$. This impedance value produces a return loss of 15 dB, representing a realistic worst-case scenario that exceeds the typical return loss specifications for our cryogenic circulators and SMA cables.
We then compare the boxcar CP-JTWPA gain profiles under different sandwiching attenuation values with this 40-$\Omega$ port impedance mismatch. Using the ideal impedance-matched transmission $|S_{21}|^{2}$ from Fig. 7 as reference, the 0-dB case (no added attenuators) shows some residual ripples from external reflections, despite the  30-dB total round-trip loss from external components alone. With 3-dB attenuators, these residual external ripples are effectively suppressed, and the gain profiles converge to levels within $\pm$0.5 dB of the 6-dB- and 10-dB-attenuator cases. Importantly, stronger attenuators do not alter the intrinsic gain profiles, as demonstrated by the nearly identical curves between 6-dB and 10-dB sandwiching attenuators. This analysis confirms that 3-dB attenuators provide sufficient isolation for the boxcar CP-JTWPA to reveal its intrinsic characteristics without introducing detectable bias in our device comparisons.